OpenStudy (anonymous):

Divide. Assume that no denominator equals zero. 10d^2/4c divided by cd^3/14 A.40c^2/d B.30c^2/d C.35/c^2d D.c^2d^5/35

OpenStudy (anonymous):

@davidmack

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

lol okay :)

OpenStudy (anonymous):

\[\frac{ \frac{ 10d^{2} }{ 4c } }{ \frac{ cd^{3} }{ 14 } } = \frac{ 10d^{2} }{ 4c }*\frac{ 14 }{ cd^{3} }\]

OpenStudy (anonymous):

now cancel like terms number with number, alphabets with alphabets

OpenStudy (anonymous):

omg thx you